OpenStudy (anonymous):

If the side od a square are langthened by 7 cm, the area becomes 144cm^2. find the langth of a side of the origonal square.

OpenStudy (callisto):

let x be the original length (x+7)^2 = 144 (x+7) = 12 or (x+7) = -12 (rejected) x=5 cm therefore, the length of the original square is 5 cm

OpenStudy (anonymous):

thank you that the answer i cam up with to

OpenStudy (callisto):

welcome :)